myininaya (myininaya):

another cal question: If f is differentiable function such that f(x) is never 0 and int(f(t),t=0..x)=[f(x)]^2 for all x, find f.

myininaya (myininaya):

\[\int\limits_{0}^{x}f(t) dt=[f(x)]^2\]

myininaya (myininaya):

that is what that one part says

myininaya (myininaya):

i need to massage my wrist i will be right back

myininaya (myininaya):

do you want me to tell you the first thing i would do?

OpenStudy (anonymous):

Yes

myininaya (myininaya):

take derivative of both sides

OpenStudy (across):

Differential equations?

myininaya (myininaya):

yes and we should be able to solve the differential equation pretty easily

myininaya (myininaya):

\[f(x)=2f(x)f'(x)\]

myininaya (myininaya):

\[f'(x)=\frac{1}{2}\]

myininaya (myininaya):

then integrate both sides and then you have f(x)

myininaya (myininaya):

this problem was too easy let me see if i find a harder one

OpenStudy (anonymous):

so f(x) would be 1/2 x +C ?

OpenStudy (across):

yep

myininaya (myininaya):

:)

OpenStudy (anonymous):

so f(t)= 1/2 t right?

myininaya (myininaya):

\[\int\limits_{0}^{x}(\frac{1}{2}t+C)dt=\frac{1}{2}*\frac{t^2}{2}+Ct+D=\frac{t^2}{4}+Ct+D\] we want this to be equal to \[(\frac{1}{2}t+C)^2=\frac{1}{4}t^2+tC+C^2\] so we need D=C^2

myininaya (myininaya):

oops just imagine all those t's are x's

OpenStudy (zarkon):

where is the D coming from?

myininaya (myininaya):

its a constant i added the constant after integrating

OpenStudy (zarkon):

you don't need it on a definite integral

myininaya (myininaya):

oops you are right

myininaya (myininaya):

imagine there is no D

myininaya (myininaya):

i think we need to take C to be zero like imaranmeah said

OpenStudy (zarkon):

then what about this ... "f(x) is never 0"

OpenStudy (anonymous):

because integral of 1/2 t is t^2 evalatuated from 0 to x so x^2 but that's not it

OpenStudy (zarkon):

what about when x is zero

myininaya (myininaya):

f(x)=1/2 *x

myininaya (myininaya):

oh then f would be 0 if x is 0 so this function is wrong

myininaya (myininaya):

ok so maybe this one is a little harder than i thought

OpenStudy (zarkon):

they probably meant to say that f(x) can't be identically zero

myininaya (myininaya):

do you have the james stewart 6th edition cal book?

OpenStudy (anonymous):

the one with big ring on it?

OpenStudy (zarkon):

yes

myininaya (myininaya):

a huge \[\int\limits_{}^{}\] symbol like this is on that cover yes imranmeah why is this book so popular?

myininaya (myininaya):

i don't about this question zarkon it says never zero

OpenStudy (zarkon):

I only have the 7th edition here at home...my 6th edition is in my office

myininaya (myininaya):

page 345 in my book who knows you might have the same question

OpenStudy (zarkon):

what section

OpenStudy (across):

I also have the 7th edition only :/

myininaya (myininaya):

problem plus after chapter 5

myininaya (myininaya):

problems plus*

OpenStudy (zarkon):

i don't see that problem

myininaya (myininaya):

number 3

OpenStudy (zarkon):

my #3 has stuff with integrals of exponential functions

myininaya (myininaya):

it sounds like they don't want f(x) to be 0 ever almost if it were an exponential function or a positive constant

myininaya (myininaya):

but exponential wouldn't work

OpenStudy (zarkon):

the question has no solution as written since \[0=\int\limits_{0}^{0}f(t) dt=[f(0)]^2\]

myininaya (myininaya):

you are right i don't think they put the answers to the problem plus in the back :(

OpenStudy (zarkon):

they dind't want a student to write \[f(x)=0\] as a solution

OpenStudy (zarkon):

didn't

OpenStudy (zarkon):

they just worded it incorrectly

myininaya (myininaya):

so maybe they should have said f(x) is not identically zero

myininaya (myininaya):

like you said

OpenStudy (zarkon):

yes...then the above solution... x/2 works

myininaya (myininaya):

maybe this is why its not in the 7th edition

OpenStudy (zarkon):

maybe ;)

myininaya (myininaya):

thats why joe hates calculus

OpenStudy (zarkon):

lol

OpenStudy (anonymous):

i hate calculus. >.>

OpenStudy (zarkon):

I love calculus :)

myininaya (myininaya):

joe why you hate cal so much?

OpenStudy (anonymous):

i love linear algebra.

OpenStudy (anonymous):

doesn't linear algebra use calc?

OpenStudy (anonymous):

i guess i should be a little more specific. What i hate in calculus is really just integration. I like everything else.

myininaya (myininaya):

so you hate integrating something like x^2

OpenStudy (anonymous):

no, linear algebra doesnt use calc, its the other way around! Differentiation is a Linear Transformation on the vector space of functions: \[\frac{d}{dx}(0) = 0\] \[\frac{d}{dx}(af(x)+g(x)) = a\frac{d}{dx}f(x)+\frac{d}{dx}g(x)\]

OpenStudy (across):

I love calculus, but I see why someone would hate it after being asked to integrate: \[\int{x^5sin(x)dx}\]Tee hee

OpenStudy (anonymous):

the same can be said for integration as well! so there are matrices that represent these transformations. Who needs calc? i'll use the matrix thank you very much :P

myininaya (myininaya):

i think they should teach remedial linear algebra before algebra i think easy general things should be thought first

OpenStudy (anonymous):

Transformation; is it like jacobian?

myininaya (myininaya):

taught* i hat english i swear

myininaya (myininaya):

hate*

myininaya (myininaya):

lol

myininaya (myininaya):

we need to "asianize" americans

OpenStudy (anonymous):

@imran yes, the jacobian matrix is also a linear transformation i believe.

OpenStudy (anonymous):

Yeah, I hated those; I forgot it during final . Had to cheat from person sitting ahead

OpenStudy (anonymous):

you do what you gotta do lol

OpenStudy (across):

How in the world can you cheat from the guy in front of you without being seen? xd

OpenStudy (anonymous):

the class is on incline. So we are higher than people sitting infront of us

OpenStudy (across):

Haha, I bet you pulled a Mr. Bean: http://www.youtube.com/watch?v=98V9cEYe6-A

OpenStudy (anonymous):

lol I've seen that, I did better because no one notice

OpenStudy (anonymous):

When I am in library , I think of that Mr.Bean scene